OpenStudy (anonymous):

Flipping a fair coin or a weighted coin. Assume that the weighted coin yields a heads with probability 0.4. You select one of the coins at random, and flip it 2 time, noting heads or tails with each flip. What is the probability that the weighted coin was selected, given that all 2 flips turned up heads____________?

OpenStudy (anonymous):

How have you been?

OpenStudy (dumbcow):

Given the coin is weighted, probability of flipping 2 heads is (.4)^2 = .16 or 4/25 Let W be weighted, F means fair coin and H mean 2 heads P(H|W) = 4/25 P(H|F) = 1/4 P(H) = (1/2)(4/25) + (1/2)(1/4) = 41/200 P(W|H) = [(1/2)(4/25)]/(41/200) = 16/41

OpenStudy (dumbcow):

i am assuming there are only 2 coins to choose from...1 weighted and 1 fair is that right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you are correct